Circular No. 6021

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) For the Hubble Space Telescope (HST) observing team, H. A. Weaver,

Space Telescope Science Institute, reports the following results from their analysis of HST WFPC-2 images and FOS spectra taken up to and including July 14: “From July 14 images of fragments G = 15 and H = 14 it appears that the inner comae are becoming extremely elongated along the direction of the train. During the period from late January through the end of June, the isophotes in the inner 0′′.7 deviated from spherical symmetry by only ∼ 15 percent (i.e., the ratio of the major and minor axes of the bestfitting ellipse was only ∼ 1.15). In the July 14 images the deviation from spherical symmetry is ∼ 90 percent. This change in morphology appears to be associated with a ‘stretching’ of the coma rather than a fragmentation of the nucleus, because the images still show sharp cores much like those observed previously by HST. These core brightnesses (defined as the sum over 3×3 pixel boxes centered on the peak pixel; 1 pixel = 0′′.046) have been decreasing monotonically with time, but the decreases since January are only in the range of ∼ 30–50 percent. This decrease in the core brightness could be due to a decrease in the dust content of the core, as opposed to a decrease in the brightness in the unresolved point source. With regard to the temporal variation in somewhat larger apertures (e.g., 0′′.3, 1′′.1 and 2′′.1; cf. IAUC 6019), our detailed analysis shows very little change for most fragments. A variation of ∼ 5–10 percent is typical, but fragment V = 2 shows a ∼ 30-percent drop from January to late June. There is a small ‘hump’ in the spatial brightness profile of fragment P_{2a} = 8b_{1} at ∼ 0′′.2 from the peak pixel, but otherwise the variation in brightness and morphology associated with this fragment has not been particularly noteworthy. No emission features are seen in a 60-min FOS spectrum of the region 222–328 nm with the 3′′.7 × 1′′.2 aperture placed near fragment G = 15. No further spectroscopic exposures of the comet will be made with HST.”

“Final” astrometric observations of most of the nuclei have been reported by R. M. West, European Southern Observatory, from his measurement of images obtained on July 14.97 and 14.98 UT by P. Dubath with the New Technology Telescope at La Silla. The final observation of nucleus C = 19 was obtained at ESO on July 14.00 UT. The final observation of nucleus B = 20 was reduced by W. Owen from an image obtained by A. and D. Monet with the U.S. Naval Observatory’s 1.5-m astrometric reflector on July 13.17. Nuclei A = 21 and D = 18 were last recorded at ESO on July 12.99 and nuclei T = 4 and U = 3 on July 7.98.

1994 July 16 Brian G. Marsden